# Beyond Face Swapping: A Diffusion-Based Digital Human Benchmark for Multimodal Deepfake Detection

### v1

@@ Abstract.
In recent years, the rapid development of deepfake technology has given rise to an emerging and serious threat to public security: diffusion model-based digital human generation. Unlike traditional face manipulation methods, such models can generate highly realistic videos with consistency through multimodal control signals. Their flexibility and covertness pose severe challenges to existing detection strategies. To bridge this gap, we introduce DigiFakeAV [^1] https://hubeiwuhanliu.github.io/DigiFakeAV.github.io/ , the first large-scale multimodal digital human forgery dataset based on diffusion models. Employing five latest digital human generation methods (Sonic, Hallo, etc.) and voice cloning method, we systematically produce a dataset comprising 60,000 videos (8.4 million frames), covering multiple nationalities, skin tones, genders, and real-world scenarios, significantly enhancing data diversity and realism. User studies show that the confusion rate between forged and real videos reaches 68%, and existing state-of-the-art (SOTA) detection models exhibit large drops in AUC values on DigiFakeAV, highlighting the challenge of the dataset. To address this problem, we further propose DigiShield, a detection baseline based on spatiotemporal and cross-modal fusion. By jointly modeling the 3D spatiotemporal features of videos and the semantic-acoustic features of audio, DigiShield achieves SOTA performance on both the DigiFakeAV and DF-TIMIT datasets. Experiments show that this method effectively identifies covert artifacts through fine-grained analysis of the temporal evolution of facial features in synthetic videos.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
@@ 1. Introduction
The rapid advancement of deepfake technology has posed unprecedented threats to digital security, giving rise to malicious applications such as political disinformation (Westerlund2019, ), synthetic pornographic content (Mirsky2021, ), and sophisticated financial fraud (yahoo2024hongkongscam, ). While existing detection benchmarks have driven progress, they remain constrained by outdated synthesis paradigms. DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be categorized into three generations. First-generation datasets (e.g., FF++ (rössler2019faceforensicslearningdetectmanipulated, )) suffer from limited scale and quality. The second generation (e.g., Celeb-DF (li2020celebdflargescalechallengingdataset, )) and third generation datasets (e.g., DFDC and FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, )) further enhanced the quantity, authenticity, and diversity of data. However, third-generation datasets primarily focus on face-swapping methods (e.g., FSGAN, Faceswap) (kaur2024hindiaudiovideodeepfakehavdfhindi, ), leading to homogenized artifacts, audiovisual mismatches, and poor generalization to emerging threats.
@@
Crucially, as diffusion models revolutionize synthetic media (chen2024echomimiclifelikeaudiodrivenportrait, ; lin2025omnihuman1rethinkingscalinguponestage, ; xu2024hallohierarchicalaudiodrivenvisual, ; cui2024hallo2longdurationhighresolutionaudiodriven, ; wang2024vexpressconditionaldropoutprogressive, ; ji2024sonicshiftingfocusglobal, ; tian2024emoemoteportraitalive, ), existing datasets fail to capture the multimodal realism and temporal coherence of modern digital human forgeries. As shown in Tab. 1, this has caused detection systems to approach performance saturation on current benchmarks while remaining fragile in real-world scenario (croitoru2024deepfakemediagenerationdetection, ). There is an urgent need for more challenging and representative datasets and benchmarks to address emerging threats.
@@
To bridge this gap, we introduce DigiFakeAV, the first large-scale multimodal benchmark for detecting diffusion-based digital human forgeries. Our dataset advances the field through three key innovations:
@@
Figure 2. Concrete examples demonstrating the diversity of DigiFakeAV, including multiple nationalities, diverse skin tones, various genders, and a range of scenarios.
@@
$\bullet$ Diffusion-Driven Synthesis: Unlike previous benchmarks relying on GANs or face-swapping, DigiFakeAV employs five state-of-the-art diffusion models ( Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), etc.) to generate 60,000 high-resolution videos (8.4M frames). These models leverage multimodal signals (such as reference images and audio inputs) to create forgeries, eliminating source-target appearance constraints while achieving photorealistic details, such as subpixel skin textures.
@@
$\bullet$ Multimodal Consistency: We ensure strict alignment between synthesized lip movements, facial expressions, and head poses with audio prosody and semantic content through speech-driven audiovisual synchronization. Compared with the baseline based on Wav2Lip, lip synchronization accuracy is significantly improved.
@@
$\bullet$ Scene-Aware Diversity: DigiFakeAV covers various real-world scenarios (news broadcasts, social media vlogs, multilingual interviews) with balanced demographic representation: 57% male/43% female, as shown in Fig. 2. We also increase the proportion of Asians and Africans, which is in stark contrast to existing datasets, reducing racial bias while enhancing the robustness of cross-cultural detection.
@@
Experimental validation reveals alarming vulnerabilities in current detectors: compared to face-swapping benchmarks, state-of-the-art models (e.g., SFIConv (10286083, )) exhibit a 43.5% drop in AUC on DigiFakeAV. Human evaluators misclassified 68% of synthetic videos as real, nearly triple the error rate of DFDC. These results underscore the urgency of developing detectors capable of identifying next-generation forgeries.
To advance research, we further propose DigiShield, a spatiotemporal audio-visual fusion framework that identifies synthetic artifacts through cross-modal inconsistency learning. By establishing the first effective baseline (AUC=80.1%), our work not only exposes critical gaps in detection research but also provides actionable insights for combating diffusion-based threats. DigiFakeAV represents a paradigm shift in deepfake benchmarking, prioritizing real-world relevance and adaptability to rapidly evolving generative technologies.
@@
$\bullet$ We introduce DigiFakeAV, the first large-scale multimodal benchmark for detecting diffusion-based digital human forgeries, comprising 60,000 high-resolution videos (8.4M frames) generated by five state-of-the-art diffusion models.
$\bullet$ Comprehensive experiments demonstrate that state-of-the-art detectors suffer a large AUC drop on DigiFakeAV, with human evaluators misclassifying 68% of synthetic videos as real, highlighting the urgency of advanced detection methods.
$\bullet$ We propose DigiShield, a spatiotemporal audiovisual fusion framework that identifies synthetic artifacts via cross-modal inconsistency learning, establishing an effective baseline for detecting next-level deepfakes.
Table 1. Performance (AUC) of state-of-the-art detection methods on existing deepfake video datasets. Results indicate that current benchmarks are no longer sufficiently challenging, motivating the creation of DigiFakeAV.
@@ 2.1. Deepfake Detection Datasets
DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be categorized into three generations. The first-generation deepfake video datasets, such as DF-TIMIT, UADFV (yang2018exposingdeepfakesusing, ), and FaceForensics++, suffer from limitations in both scale and quantity—for example, UADFV contains only 98 videos. These early benchmark datasets laid the foundation for preliminary detection research, but are no longer applicable. The second-generation datasets, including Celeb-DF and DeeperForensics, have improved in scale and quality compared to the first generation, encompassing more forgery techniques and larger sample sizes. However, they still lack diversity in forgery types and multimodal synthetic content. Represented by DFDC and FakeAVCeleb, third-generation datasets include tens of thousands of samples and introduce audio synthesis technologies. DFDC contains synthetic but asynchronous audio, while FakeAVCeleb uses SV2TTS (jia2019transferlearningspeakerverification, ) and Wav2Lip for speech cloning and video reconstruction. Although speech synthesis technologies have been updated, other forgery techniques remain largely unchanged from second-generation datasets.
With the advancement of generative models, diffusion-based forgery has transcended facial regions and lip synchronization to involve multimodal digital human generation, achieving unprecedented levels of realism and audiovisual coherence. These technologies pose severe threats to information security, social trust, and media authenticity. To address this challenge, we created DigiFakeAV, a diffusion-based dataset featuring large-scale, high-quality, multimodal synthetic videos across diverse scenarios, providing a more comprehensive benchmark for robust deepfake detection.
@@ 2.2. Deepfake Video Generation
The progress of generative models has significantly advanced the evolution of deepfake videos. Early research in this field primarily relied on traditional GAN-based models for face swapping and reconstruction. For example, various multimodal forged video datasets employ Wav2Lip (Prajwal_2020, ), a method that utilizes a discriminator mechanism to improve the accuracy of audiovisual synchronization. Gao et al. (gao2023high, ) propose a face reenactment GAN focused on generating head movements synchronized with speech. However, GAN-based methods still face challenges such as unnatural head movements, facial distortions, and inaccurate identity retention. Additionally, training GANs requires a delicate balance between generators and discriminators, often leading to instability and sensitivity to data distributions.
In recent years, diffusion models have revolutionized the generation of videos based on portraits and digital humans. These methods take audio and reference images as input prompts, driving consistent and expressive animations through diffusion processes. For example, Animate Someone (hu2024animateanyoneconsistentcontrollable, ) can generate expressive videos with minimal control inputs. Based on this, EMO (tian2024emoemoteportraitalive, ) introduces a frame encoding module and improves temporal control to achieve coherent audio-driven animations. V-Express (wang2024vexpressconditionaldropoutprogressive, ) enhances audiovisual alignment through a hierarchical architecture, enabling precise synchronization between speech, lip movements, facial expressions, and head poses. Hallo (xu2024hallohierarchicalaudiodrivenvisual, ) proposes a hierarchical synthesis framework that effectively handles lip synchronization, expression, and pose alignment in portrait animation. In contrast, Sonic (ji2024sonicshiftingfocusglobal, ) adopts a global audio-aware design, achieving efficient long-video generation via time-shifted fusion without motion frames, while reducing computational overhead.
Synthetic videos generated by these diffusion-based methods are increasingly indistinguishable from real ones, marking the rise of next-level digital human-based forgery technologies and the obsolescence of current deepfake detection datasets. Therefore, there is an urgent need to establish new benchmarks based on diffusion models to drive corresponding advancements in detection and ensure their robustness against emerging forgery methods.
3. DigiFakeAV
In this section, we describe the pipeline of dataset construction, including data collection, data synthesis, and the composition of the final dataset, as shown in Fig. 3. In the core subsection of data synthesis, we discuss the specific generation methods.
@@ 3.1. Data Collection
The real videos used in this study are sourced from the HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) datasets. We implement a dedicated video processing pipeline to split all MP4, AVI, and MKV files in the original directories into shorter clips, which are then cropped and re-encoded. Specifically, for each video file, we first calculate its total duration and split it into segments of 3–5 seconds each.
Subsequently, we use FFmpeg to process each clip, including resizing to 512×512 pixels, copying the audio stream without re-encoding, and setting the bitrate to 3000 kbps to balance processing speed and quality. Through the above steps, we obtain nearly 40,000 clean video clips from the HDTF and CelebV-HQ datasets for further analysis.
@@ 3.2. Data Synthesis
We produce the DigiFakeAV dataset using state-of-the-art deepfake and audio synthesis models. Specifically, we employ five recent portrait animation video generation methods and one advanced audio synthesis method. As summarized in Tab. 2, compared with previous deepfake generation methods, our choices exhibit significant differences and advantages.
Synthesis Pipeline. Both methods generate fake video and fake audio samples through three stages: (1) Generation conditioning: The video generation model is driven by real audio (from real video–real audio clips) or synthetic audio (e.g., TTS), paired with static portraits or video frames as visual inputs. (2) Diffusion sampling: Video frames are synthesized via iterative denoising guided by audio and facial priors. Clips are 5–10 seconds long, meeting the dataset’s resolution (512×512) and bitrate (3000 kbps) requirements. (3) Quality control: Samples are filtered using perceptual metrics (e.g., FID, Sync-C) to eliminate artifacts such as lip-sync errors or identity drift. Realistic noise and compression are added to simulate real-world distribution shifts.
Key Advantages. (1) Higher Realism: Diffusion models can generate fine-grained facial details (e.g., skin textures, eye movements) and maintain long-term temporal consistency, outperforming GAN-based methods. (2) Audiovisual Consistency: Sonic and Hallo achieve strong lip-sync accuracy and emotional consistency, challenging unimodal detection. (3) Scalability: Both frameworks support batch processing and domain-specific fine-tuning (e.g., celebrity voice dubbing), enabling rapid dataset expansion.
By integrating these methods, DigiFakeAV provides rich FV-FA samples that closely resemble real-world forgery scenarios, thus supporting the development of robust audiovisual deepfake detection systems.

### v2

@@ Abstract.
In recent years, the rapid development of deepfake technology has given rise to an emerging and serious threat to public security: diffusion-based digital human generation. Unlike traditional face manipulation methods, such models can generate highly realistic videos with consistency through multimodal control signals. Their flexibility and covertness pose severe challenges to existing detection strategies. To bridge this gap, we introduce DigiFakeAV [^1] https://hubeiwuhanliu.github.io/DigiFakeAV.github.io/, the new large-scale multimodal digital human forgery dataset based on diffusion models. Employing five of the latest digital human generation methods and the voice cloning methods, we systematically produce a dataset comprising 60,000 videos (8.4 million frames), covering multiple nationalities, skin tones, genders, and real-world scenarios, significantly enhancing data diversity and realism. User studies demonstrate that participants misclassify forged videos as real in 68% of tests, and existing detection models exhibit a large drop in performance on DigiFakeAV, highlighting the challenge of the dataset. To address this problem, we propose DigiShield, an effective detection baseline based on spatiotemporal and cross-modal fusion. By jointly modeling the 3D spatiotemporal features of videos and the semantic-acoustic features of audio, DigiShield achieves state-of-the-art (SOTA) performance on the DigiFakeAV and shows strong generalization on other datasets.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference’33; October 2025; Dublin, Irelandbooktitle: Conference’33, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
The rapid advancement of deepfake technology has posed unprecedented threats to digital security, giving rise to malicious applications such as political disinformation (Westerlund2019, ), synthetic pornographic content (Mirsky2021, ), and sophisticated financial fraud (yahoo2024hongkongscam, ). While existing detection benchmarks have driven progress, they remain constrained by outdated synthesis paradigms. The existing deepfake detection datasets can be roughly categorized into three generations (dolhansky2020deepfakedetectionchallengedfdc, ). First-generation datasets (e.g., FF++ (rössler2019faceforensicslearningdetectmanipulated, )) suffer from limited scale and quality. The second generation (e.g., Celeb-DF (li2020celebdflargescalechallengingdataset, )) and third generation datasets (e.g., DFDC (dolhansky2020deepfakedetectionchallengedfdc, )) further enhanced the quantity, authenticity, and diversity of data. However, the latest third-generation datasets primarily focus on face-swapping methods (e.g., FSGAN, Faceswap) (kaur2024hindiaudiovideodeepfakehavdfhindi, ), leading to homogenized artifacts, audiovisual mismatches, and poor generalization to emerging threats.
@@
Crucially, as diffusion models revolutionize synthetic media (chen2024echomimiclifelikeaudiodrivenportrait, ; lin2025omnihuman1rethinkingscalinguponestage, ; xu2024hallohierarchicalaudiodrivenvisual, ; cui2024hallo2longdurationhighresolutionaudiodriven, ; wang2024vexpressconditionaldropoutprogressive, ; ji2024sonicshiftingfocusglobal, ; tian2024emoemoteportraitalive, ), existing datasets fail to capture the multimodal realism and temporal coherence of modern digital human forgeries. As shown in Tab. 1, this has caused detection systems to approach performance saturation on current benchmarks while remaining fragile in real-world scenarios (croitoru2024deepfakemediagenerationdetection, ). There is an urgent need for more challenging and representative datasets and benchmarks to address emerging threats.
@@
To bridge this gap, we introduce DigiFakeAV, the new large-scale multimodal benchmark for detecting diffusion-based digital human forgeries. Our dataset advances the field through three key innovations:
@@
Figure 2. Examples demonstrating the diversity of DigiFakeAV, including multiple nationalities, diverse skin tones, various genders, and a range of scenarios.
[…]
Examples demonstrating the diversity of DigiFakeAV
@@
$\bullet$ Diffusion-Driven Synthesis: unlike previous benchmarks relying on GANs or face-swapping, DigiFakeAV employs five state-of-the-art diffusion models ( Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), etc.) to generate 60,000 high-resolution videos (8.4 million frames). These models leverage multimodal signals (such as reference images and audio inputs) to create forgeries, eliminating source-target appearance constraints while achieving photorealistic details, such as subpixel skin textures.
@@
$\bullet$ Multimodal Consistency: we ensure strict alignment between synthesized lip movements, facial expressions, and head poses with audio prosody and semantic content through speech-driven audiovisual synchronization. Lip synchronization is significantly improved compared with the previous datasets based on Wav2Lip (Prajwal_2020, ).
@@
$\bullet$ Scene-Aware Diversity: DigiFakeAV covers various real-world scenarios (news broadcasts, social media vlogs, multilingual interviews) with balanced demographic representation: 57% male/43% female, as shown in Fig. 2. We also increase the proportion of Asians and Africans, reducing bias while enhancing the robustness of cross-cultural detection.
@@
Experiments reveal alarming vulnerabilities in current detectors: compared to face-swapping benchmarks, SOTA models experience a dramatic drop in Area Under the Curve (AUC) on DigiFakeAV. For example, SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, ) exhibits a 43.5% decrease. Human evaluators misclassified 68% of synthetic videos as real, nearly triple the error rate of DFDC (dolhansky2020deepfakedetectionchallengedfdc, ). This suggests that we must design detectors that can identify next-generation forgeries.
To address the challenge of diffusion-based digital human forgeries, we propose DigiShield, a new multimodal detection framework. The framework explicitly captures dynamic inconsistencies between video and audio through cross-modal attention mechanisms and spatiotemporal feature extraction networks, enabling effective identification of subtle forgery traces generated by diffusion models. Experiments show that DigiShield outperforms existing methods on the DigiFakeAV dataset, validating its effectiveness in next-generation deepfake detection and providing scalable framework design insights for future research.
@@
$\bullet$ We introduce DigiFakeAV, the new large-scale multimodal benchmark for detecting diffusion-based digital human forgeries, comprising 60,000 high-resolution videos (8.4 million frames) generated by SOTA diffusion models.
$\bullet$ Comprehensive experiments demonstrate that SOTA detectors suffer a large performance drop on DigiFakeAV, with human evaluators misclassifying 68% of synthetic videos as real, highlighting the urgency of advanced detection methods.
$\bullet$ We propose DigiShield, a spatiotemporal audiovisual fusion framework that identifies synthetic artifacts via cross-modal inconsistency learning, establishing a new baseline for detecting next-level deepfakes.
Table 1. Performance (AUC) of SOTA detection methods on existing deepfake video datasets. Results indicate that current benchmarks are constrained by outdated synthesis paradigms.
@@ 2.1. Deepfake Detection Datasets
DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be roughly categorized into three generations. The first-generation deepfake video datasets, such as DF-TIMIT (korshunov2018deepfakesnewthreatface, ), UADFV (yang2018exposingdeepfakesusing, ), and FF++  (rössler2019faceforensicslearningdetectmanipulated, ), suffer from limitations in both scale and quantity. For example, UADFV contains only 98 videos. These early benchmark datasets laid the foundation for initial detection research, but are no longer applicable. The second-generation datasets, including Celeb-DF (li2020celebdflargescalechallengingdataset, ) and DeeperForensics (jiang2020deeperforensics10largescaledatasetrealworld, ), have improved in scale and quality compared to the first generation, encompassing more forgery techniques and larger sample sizes. However, they still lack diversity in forgery types and multimodal synthetic content. Represented by DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) and FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ), third-generation datasets include tens of thousands of samples and introduce audio synthesis technologies. DFDC contains synthetic but asynchronous audio, while FakeAVCeleb uses SV2TTS (jia2019transferlearningspeakerverification, ) and Wav2Lip (Prajwal_2020, ) for speech cloning and video reconstruction. Except for speech synthesis technologies, other forgery techniques remain unchanged from the second generation.
With advancements in generative models, diffusion-based forgery (wei2024aniportraitaudiodrivensynthesisphotorealistic, ) now enables highly realistic and coherent multimodal digital human generation beyond facial and lip synthesis. To address the detection challenges this poses, we introduce DigiFakeAV, a high-quality, large-scale diffusion-based dataset of multimodal synthetic videos spanning diverse scenarios, offering a comprehensive benchmark for robust deepfake detection.
@@ 2.2. Deepfake Video Generation
The progress of generative models has significantly advanced the evolution of deepfake videos. Early research primarily relies on traditional GAN-based models for face swapping and reconstruction. For example, various multimodal forged video datasets employ Wav2Lip (Prajwal_2020, ), utilizing a discriminator mechanism to improve audiovisual synchronization. Gao et al. (gao2023high, ) propose a face reenactment GAN focused on generating head movements synchronized with speech. However, GAN-based methods still face challenges such as unnatural head movements, facial distortions, and inaccurate identity retention. Additionally, training GANs often leads to instability and sensitivity to data distributions.
In recent years, diffusion models have revolutionized the generation of digital human videos. These methods take audio and reference images as input prompts, driving consistent and expressive animations through diffusion processes. For example, Animate Someone  (hu2024animateanyoneconsistentcontrollable, ) can generate expressive videos with minimal control inputs. Based on this, EMO (tian2024emoemoteportraitalive, ) introduces a frame encoding module and improves temporal control to achieve coherent audio-driven animations. V-Express (wang2024vexpressconditionaldropoutprogressive, ) enhances audiovisual alignment through a hierarchical architecture. Hallo (xu2024hallohierarchicalaudiodrivenvisual, ) proposes a hierarchical synthesis framework that effectively handles lip synchronization. Moreover, Sonic (ji2024sonicshiftingfocusglobal, ) adopts a global audio-aware design, achieving efficient long-video generation via time-shifted fusion without motion frames.
Synthetic videos generated by these diffusion-based methods are indistinguishable from real ones, marking the rise of next-level forgery technologies and the obsolescence of current deepfake detection datasets. Therefore, there is an urgent need to establish new benchmarks based on diffusion models to drive corresponding advancements in detection and ensure their robustness against emerging forgery methods.
3. DigiFakeAV: A Next Generation Deepfake Benchmark
In this section, we describe the pipeline of dataset construction, including data collection, data synthesis, and the composition of the final dataset, as shown in Fig. 3. In the core subsection of data synthesis, we discuss the specific generation methods [^2] For further details of our dataset, please refer to the Supplementary Materials (Appendix A), at https://github.com/DigiFakeAV/DigiFakeAV-v1/blob/main/supplementary/Supplementary.pdf .
@@ 3.1. Data Collection
The real videos of our dataset are sourced from the HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) datasets. We implement a dedicated video processing pipeline to split all MP4, AVI, and MKV files in the original directories into shorter clips, which are then cropped and re-encoded. Specifically, for each file, we first calculate its total duration and split it into segments of 3–5 seconds each. Subsequently, we use FFmpeg to process each clip, including resizing to 512×512 pixels, copying the audio stream without re-encoding, and setting the bitrate to 3000 kbps to balance processing speed and quality.
Through the above steps, we obtain nearly 40,000 clean video clips from the HDTF and CelebV-HQ for further analysis.
@@ 3.2. Data Synthesis
We produce the DigiFakeAV dataset using state-of-the-art (SOTA) deepfake and audio synthesis models. Specifically, we employ five recent diffusion-based digital human video generation methods (Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), EchoMimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, )) and one advanced audio synthesis method (CosyVoice 2 (du2024cosyvoice2scalablestreaming, )). As summarized in Tab. 2, compared with previous deepfake methods, our choices exhibit significant advantages.
Synthesis Pipeline. Both methods generate fake video and fake audio samples through three stages: (1) Generation conditioning: the video generation model is driven by real audio (from real video–real audio clips) or synthetic audio (e.g., TTS), paired with static portraits or video frames as visual inputs. (2) Diffusion sampling: video frames are synthesized via iterative denoising guided by audio and facial priors. Clip lengths range from 5 to 10 seconds, meeting the dataset’s resolution (512×512) and bitrate (3000 kbps) requirements. (3) Quality control: samples are filtered using perceptual metrics (e.g., FID, Sync-C) to eliminate artifacts such as lip-sync errors or identity drift [^3] Generation quality Standards are in Appendix A.3. Realistic noise and compression are added to simulate real-world distribution shifts.
Key Advantages. (1) Higher Realism: diffusion models can generate fine-grained facial details (skin textures, eye movements) and maintain long-term temporal consistency, outperforming GAN-based methods. (2) Audiovisual Consistency: Sonic and Hallo achieve strong lip-sync accuracy and emotional consistency, challenging unimodal detection. (3) Scalability: both frameworks support batch processing and domain-specific fine-tuning (e.g., celebrity voice dubbing), enabling rapid dataset expansion.
By integrating these methods, DigiFakeAV provides rich fake video and fake audio samples that closely resemble real-world forgery scenarios, thus supporting the development of robust deepfake detection systems.